OpenStudy (anonymous):

How do you find the slope-intercept form of? -1/4x - y - 1/2x = 5?

OpenStudy (ryan123345):

add the like terms.. add the fractions.

OpenStudy (anonymous):

combine like terms of 1/2x and -1/4x. That is the slope. Next, slope intercept form is y=mx+b (b is the y intercept)

OpenStudy (anonymous):

subtract x's from left side and place it on the right side

OpenStudy (anonymous):

then divide by -1 since you will have -y=mx+b

OpenStudy (anonymous):

Is it \[y = -\frac{ .25 }{ 1 }x+-5\]?

OpenStudy (anonymous):

How did you get .25? This is what i got: \[y = -3/4x - 5\]

OpenStudy (anonymous):

combining -1/4 and -1/2 = -3/4